OpenStudy (anonymous):

Please help me solve this type of problem and show steps: (6-3i)+(5+4i) i = imaginary #

jimthompson5910 (jim_thompson5910):

(6-3i)+(5+4i) 6-3i+5+4i (6+5)+(-3i+4i) 11+i So (6-3i)+(5+4i) = 11+i

OpenStudy (anonymous):

just add the real parts and imaginary parts ( containing i) = 11 + i

OpenStudy (anonymous):

Thanks :D My math teacher told us to do distributive property... ????? What?! LOL, i dont like him, (:

jimthompson5910 (jim_thompson5910):

He's basically saying that (6-3i)+(5+4i) is really 1(6-3i)+1(5+4i) which distributes to 6-3i+5+4i